OpenStudy (anonymous):

does f(2) mean the same thing as f(x)=2?

OpenStudy (anonymous):

i think so

hero (hero):

Um NO!

OpenStudy (anonymous):

jay123 i got 9 7/6 ... how does it change to 10 1/6

hero (hero):

f(2) means evaluate the function at x = 2

hero (hero):

For example Suppose f(x) = x^2 and we had to evaluate f(2) Then we would take f(x) = x^2 and replace everywhere we see x with 2, then evaluate as follows.... f(2) = 2^2 f(2) = 4

OpenStudy (anonymous):

f(2) , here x=2 which is an element in the domain wheras f(x)=2 , that is the y value which is an element from the range.